OpenStudy (anonymous):

short but sweet question:

OpenStudy (anonymous):

find an integer n such that: \[n^2 < 33127 < (n+1)^2\] find a small integer m such that: \[(n+m)^2 - 33127 \] is a perfect square hence write 33127 in the form\[ pq\] where p and q are integers greater than 1.

OpenStudy (experimentx):

n = floor(sqrt(33127))

OpenStudy (experimentx):

(m+n)^2 = m^2 + 2mn + n^2 - 33127 => 33127 = 4 mn

OpenStudy (experimentx):

Oo ... something went wrong LOL

OpenStudy (anonymous):

its a non calculator, i am very interested in seeing if there is any nice way of getting the first part, when i did it i used a little trick, but its not much

OpenStudy (experimentx):

7 am and haven't slept ... i'll book mark this thread. LOL

OpenStudy (anonymous):

haha fair enough, i am also very tired, heading to bed now. night!

OpenStudy (experimentx):

GN

OpenStudy (anonymous):

for first part take sqrt of 33127 = 182.0082... so just try n=182 first which satisfy our euation so n = 182

OpenStudy (anonymous):

yeah, although i wouldn't have a calculator to sqrt, i suppose i could use a numerical method i saw that 180^2 = 32400 33127 - 32400 = 727 so we are looking for t such that 360t + t^2 = 727 and t = 2 gets us close so n = 182

OpenStudy (experimentx):

thats cheating ... LOL

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

try the other parts if you want

OpenStudy (anonymous):

the only clever trick in this question really is the pq bit

OpenStudy (anonymous):

i felt so stupid when i realised what to do, i had been sitting there for a while lol